# Tuning Stability of AB3-Type Alloys by Suppressing Magnetism

### v1

@@ Abstract
Keywords: Metal Alloy, Coherent Potential Approximation, Stability, Magnetic Properties, Hydrogen Storage
@@ 3.1 ACo3 Systems: Strong Magnetism Limits Stability
The formation energy as a function of total magnetic moment for CaxYyMg1-x-yCo3 alloys, evaluated using three reference structures and the averaged values from Eq. 2, is shown in Figure 3(d). In the CaCo3 and YCo3 structures, which possess relatively large unit-cell volumes, a clear one-to-one correlation between total magnetic moment and formation energy is observed: the formation energy systematically increases with increasing magnetization. This monotonic behavior indicates that magnetism is the dominant factor governing alloy stability under these structural conditions. In contrast, the MgCo3 structure, with its much smaller equilibrium volume, exhibits no such one-to-one correspondence–compositions with similar total magnetic moments show distinct formation energies. This breakdown of the correlation suggests that in the compact MgCo3 lattice, magnetism couples more strongly with lattice volume and electronic degrees of freedom as the strain effect, thereby weakening the direct magnetism–stability dependence. From the magnetization ternary diagrams in Figure 2(d–f), compositions with low total magnetic moments (left-hand side) correspond to Y-rich regions near YCo3, whereas those with high magnetic moments (right-hand side) correspond to Mg-rich regions near MgCo3. Among the three end-member compounds, each structure attains its lowest formation energy when evaluated using its own lattice parameters. The average formation energy versus average total magnetic moment derived from Eq. 2 smoothly approaches the corresponding stoichiometric structures as the A-site composition approaches their respective limits: starting from YCo3 with the lowest formation energy, crossing the CaCo3 structure, and finally reaching MgCo3 with the highest magnetization.
@@
Monte Carlo simulations yield more realistic estimates of the magnetic transition temperatures. As shown in Figure 5(a), the temperature dependence of magnetization and susceptibility for CaCo3, YCo3, and MgCo3 clearly reveals ferromagnetic-to-paramagnetic transitions, with Curie temperatures identified from the susceptibility maxima: 602 K (CaCo3), 474 K (YCo3), and 697 K (MgCo3). These values, while lower than the mean-field estimates, preserve the same trend of suppressed ordering temperature in YCo3. The magnetic energy profiles in Figure 5(b) further highlight this behavior. All compounds exhibit large negative magnetic energies at 0 K, with MgCo3 showing the most negative value, consistent with its highest saturation magnetization. The steepest energy variations occur at the Curie temperatures, reflecting the peak in magnetic heat capacity. Together, these results demonstrate that strong ferromagnetism, manifested in high magnetic moments and elevated Curie temperatures, correlates with reduced thermodynamic stability in Co-based AB3 alloys. Notably, Mg incorporation, while favorable for achieving high hydrogen storage density, amplifies ferromagnetic interactions and thereby destabilizes the lattice. This intrinsic trade-off underscores the need to explore alternative B-site elements with weaker magnetic interactions to balance hydrogen storage capacity and structural stability.
@@ 3.2 ANi3 Systems: Reduced Magnetism Enables Enhanced Stability
The averaged formation energy and total magnetic moment diagrams for CaxYyMg1-x-yNi3 are presented in Figure 9(a, b), with the corresponding theoretical gravimetric hydrogen densities from Eq. 3 shown in Figure 9(c). The universally negative formation energies mark a significant improvement over Co-based systems, as compositions with high hydrogen storage potential are now thermodynamically accessible. The hydrogen density distribution closely mirrors that of the Co system, with maximum values near the Mg-rich corner reaching $\sim$3.40 wt% for pure MgNi3H7. Importantly, the experimentally known composition CaMg2Ni9 (Ca0.33Mg0.67Ni3) achieves both high hydrogen storage density ($\sim$3.32 wt%) and sufficient thermodynamic stability11, 19. Most significantly, our calculations identify the previously unexplored Mg-rich corner of the ternary diagram as a promising compositional region that combines high theoretical hydrogen capacity with robust stability. This finding extends beyond known experimental compositions, highlighting new synthetic targets for high-performance hydrogen storage materials. The key enabling factor is the weak magnetism in Ni-based systems. Unlike Co-based alloys, where strong ferromagnetism imposes a fundamental stability barrier for high-capacity compositions, Ni-based AB3 alloys maintain thermodynamic stability across the entire compositional range while preserving comparable hydrogen storage densities. This demonstrates that suppression of magnetic interactions is critical for reconciling high hydrogen capacity with structural stability in AB3-type hydrides.

### v2

@@ 2 Methods
The lattice volumes of CaB3, YB3, and MgB3 (B = Co, Ni) differ significantly, with MgB3 being much smaller than CaB3 and YB3 due to small atomic radius of Mg. Although the coherent potential approximation (CPA) allows estimation of formation energies for (Ca, Y, Mg)B3 alloys with disorder effects at moderate computational cost, it cannot optimize lattice parameters, limiting its ability to capture lattice volume-dependent effects. Our averaging scheme assumes that lattice volume changes primarily arise from the A-site element. As a result, the formation energy of A-rich corners (A: Ca, Y, Mg) approaches the formation energy of the corresponding stoichiometric AB3 structure with fully relaxed by using VASP. This approach captures lattice volume-dependent effects more accurately than a simple equal-weight average, which will not approach the formation energy of the corresponding stoichiometric AB3 structure. Moreover, it includes the formation energies of two other structures, which are not the ground state structure of the corresponding stoichiometric composition.
@@ 3.1 ACo3 Systems: Strong Magnetism Limits Stability
The formation energy as a function of total magnetic moment for CaxYyMg1-x-yCo3 alloys, evaluated using three reference structures and the averaged values from Eq. 2, is shown in Figure 3(d). In the CaCo3 and YCo3 structures, which possess relatively large unit-cell volumes, a clear one-to-one correlation between total magnetic moment and formation energy is observed: the formation energy systematically increases with increasing magnetization. This monotonic behavior indicates that magnetism is the dominant factor governing alloy stability under these structural conditions. In contrast, the MgCo3 structure, with its much smaller equilibrium volume, exhibits no such one-to-one correspondence–compositions with similar total magnetic moments show distinct formation energies. This breakdown of the correlation suggests that in the compact MgCo3 lattice, magnetism couples more strongly with lattice volume and electronic degrees of freedom as the strain effect, thereby weakening the direct magnetism–stability dependence. In CaCo3 and YCo3, the lattice volumes are sufficiently large to sustain local magnetic moments, resulting in an almost linear dependence of formation energy on total magnetic moments. In MgCo3, the smaller lattice volume introduces strain effects that perturb the relationship between formation energy and magnetization, leading to a non-linear behavior. This highlights the importance of lattice-volume effects on the coupling between magnetism and formation energy.
From the magnetization ternary diagrams in Figure 2(d–f), compositions with low total magnetic moments (left-hand side) correspond to Y-rich regions near YCo3, whereas those with high magnetic moments (right-hand side) correspond to Mg-rich regions near MgCo3. Among the three end-member compounds, each structure attains its lowest formation energy when evaluated using its own lattice parameters. The average formation energy versus average total magnetic moment derived from Eq. 2 smoothly approaches the corresponding stoichiometric structures as the A-site composition approaches their respective limits: starting from YCo3 with the lowest formation energy, crossing the CaCo3 structure, and finally reaching MgCo3 with the highest magnetization.
@@
Monte Carlo simulations yield more realistic estimates of the magnetic transition temperatures. As shown in Figure 5(a), the temperature dependence of magnetization and susceptibility for CaCo3, YCo3, and MgCo3 clearly reveals ferromagnetic-to-paramagnetic transitions, with Curie temperatures identified from the susceptibility maxima: 602 K (CaCo3), 474 K (YCo3), and 697 K (MgCo3). These values, while lower than the mean-field estimates, preserve the same trend of suppressed ordering temperature in YCo3. The Curie temperature of YCo3 in previous experimental work is 301 K, which is lower than the value of Monte Carlo simulations30. The magnetic energy profiles in Figure 5(b) further highlight this behavior. All compounds exhibit large negative magnetic energies at 0 K, with MgCo3 showing the most negative value, consistent with its highest saturation magnetization. The steepest energy variations occur at the Curie temperatures, reflecting the peak in magnetic heat capacity. Together, these results demonstrate that strong ferromagnetism, manifested in high magnetic moments and elevated Curie temperatures, correlates with reduced thermodynamic stability in Co-based AB3 alloys. Notably, Mg incorporation, while favorable for achieving high hydrogen storage density, amplifies ferromagnetic interactions and thereby destabilizes the lattice. This intrinsic trade-off underscores the need to explore alternative B-site elements with weaker magnetic interactions to balance hydrogen storage capacity and structural stability.
@@ 3.2 ANi3 Systems: Reduced Magnetism Enables Enhanced Stability
The Liechtenstein formalism is generally suitable for estimating exchange parameters ($J_{ij}$) in itinerant magnets, such as Ni-based alloys. However, limitations arise from higher-order or non-Heisenberg interactions. For higher-order contributions, the magnetic force theorem remains valid but requires more complex multipolar rotations. Overall, these considerations provide a pathway to improve the accuracy of $J_{ij}$ while retaining the computational efficiency of the formalism.
@@
In YB3 and CaB3 (B = Co, Ni) crystal structures, the lattice volume is sufficient to sustain local magnetic moments on Co and Ni sites. This results in a near-linear dependence of formation energy on total magnetic moments, since the magnetic contribution to formation energy correlates directly with composition changes. In MgCo3, however, the smaller lattice volume introduces additional strain effects, which complicate the dependence of formation energy on total magnetic moments. Despite this, local magnetic moments of Co persist even in Ca- or Y-rich regions, as shown in Figure 2(f). In contrast, Ni exhibits weaker magnetism than Co. In the MgNi3 structure, local magnetic moments of Ni can be quenched in Ca- or Y-rich regions due to small lattice volumes, leading to a much stronger lattice-volume effect compared with Co-based alloys, as shown in Figure 8(f).
@@
The averaged formation energy and total magnetic moment diagrams for CaxYyMg1-x-yNi3 are presented in Figure 9(a, b), with the corresponding theoretical gravimetric hydrogen densities from Eq. 3 shown in Figure 9(c). The universally negative formation energies mark a significant improvement over Co-based systems, as compositions with high hydrogen storage potential are now thermodynamically accessible. The hydrogen density distribution closely mirrors that of the Co system, with maximum values near the Mg-rich corner reaching $\sim$3.40 wt% for pure MgNi3H7. Importantly, the experimentally known composition CaMg2Ni9 (Ca0.33Mg0.67Ni3) achieves both high hydrogen storage density ($\sim$3.32 wt%) and sufficient thermodynamic stability11, 19. This indicates a moderate improvement in hydrogen storage capacity when using MgNi3 compared with CaMg2Ni9. Most significantly, our calculations identify the previously unexplored Mg-rich corner of the ternary diagram as a promising compositional region that combines high theoretical hydrogen capacity with robust stability. This finding extends beyond known experimental compositions, highlighting new synthetic targets for high-performance hydrogen storage materials. The key enabling factor is the weak magnetism in Ni-based systems. Unlike Co-based alloys, where strong ferromagnetism imposes a fundamental stability barrier for high-capacity compositions, Ni-based AB3 alloys maintain thermodynamic stability across the entire compositional range while preserving comparable hydrogen storage densities. This demonstrates that suppression of magnetic interactions is critical for reconciling high hydrogen capacity with structural stability in AB3-type hydrides.
Although our calculations indicate that the Mg-rich region is thermodynamically stable, some experimental challenges may arise. Mg with high volatility requires precise stoichiometry control, and minor secondary phases may form if synthesis conditions are not optimized. Additionally, Mg-containing alloys may need careful surface treatment to ensure efficient hydrogen activation. These are typical experimental factors and do not represent fundamental limitations. Overall, the predicted stability and hydrogen storage capacity of Mg-rich Ni-based alloys remain robust, with only routine synthesis considerations expected.
@@
Since AB3 alloys can be viewed as composed of two-thirds AB2 and one-third AB5 building blocks, our computational scheme is likely to perform well for alloys with similar crystal structures. However, challenges may arise when applying this scheme to materials containing rare-earth elements, such as SmCo5 or LaNi5. In these cases, strong correlation effects of localized $f$-electrons are difficult to capture accurately within standard DFT, which could affect the predicted formation energies and magnetic properties.
In (Ca, Y, Mg)Co3 alloys, formation energy correlates with the total magnetic moment, which in turn reflects chemical composition. Mg-rich alloys exhibit high total magnetic moments and correspondingly higher formation energies (reduced stability), whereas Y-rich alloys suppress magnetism (low-spin state of Co and non-magnetic of Ni), which reduces the formation energy and enhances stability. Thus, magnetism and formation energy are strongly coupled, and the magnetic behavior, represented by the total magnetization, can serve as an indicator of alloy stability through its correlation with formation energy. The mechanism controlling the formation energy originates from the A-site composition (Ca, Y, or Mg), which strongly couples to the magnetic strength, such as the total magnetic moment.
On the other hand, the reason why the formation energy is reduced by replacing Co with Ni might come from several reasons, such as the electronic structure, $d$-band filling, and weaker magnetic energy penalty of Ni-based alloys. Ca, Y, and Mg donate electrons to form an intermetallic compound with Co and Ni. In the Co case, an unfilled d-band is less energetically favorable with charge transfer because more antibonding states will be occupied with high formation energy. Meanwhile, the nearly filled d-band of Ni becomes fully filled, which strengthens the bonding of the intermetallic compound, leading to lower formation energy. In the Co case, the material often pays an energy cost to maintain a high magnetic moment or spin polarization. Meanwhile, the Ni-based alloy has a smaller magnetic moment with a smaller energy penalty for spin polarization. Moreover, many Ni-based intermetallic alloys, such as LaNi5, and CeNi5 are very stable for hydrogen storage, while the Co-based intermetallic alloy is more rare due to high formation energy and magnetic instability.
@@ 4 Summary
Supporting Information
Derivation of H/M expressions for AB2, AB3, and AB5 alloys; rescaling formula for formation energies; lattice parameters for all stoichiometric AB3 compositions (A = Ca, Y, Mg; B = Co, Ni).